Please help me solve this problem: 1/2 + (2/3 + 3/4) - (4/5 * 5/6)=?
This is as far as I got:
2/3= (2/3)(4/4)= 8/12
3/4= (3/4)(3/3)= 9/12
?= 17/12

Answer by jim_thompson5910(35256)

Good, so 2/3 + 3/4 = 17/12 (the terms that change are in blue and change to what's shown in green)

1/2 + (2/3 + 3/4) - (4/5 * 5/6)

1/2 + 17/12 - (4/5 * 5/6)

Now keep going...

1/2 + 17/12 - 20/30

1/2 + 17/12 - 2/3

6/12 + 17/12 - 8/12

(6+17-8)/12

15/12

5/4

So 1/2 + (2/3 + 3/4) - (4/5 * 5/6) = 5/4